# AortaDiff: A Unified Multitask Diffusion Framework for Contrast-Free AAA Imaging

### v1

@@ Abstract
While contrast-enhanced CT (CECT) is standard for assessing abdominal aortic aneurysms (AAA) , the required iodinated contrast agents pose significant risks, including nephrotoxicity, patient allergies, and environmental harm. To reduce contrast agent use, recent deep learning methods have focused on generating synthetic CECT from non-contrast CT (NCCT) scans. However, most adopt a multi-stage pipeline that first generates images and then performs segmentation, which leads to error accumulation and fails to leverage shared semantic and anatomical structures. To address this, we propose a unified deep learning framework that generates synthetic CECT images from NCCT scans while simultaneously segmenting the aortic lumen and thrombus. Our approach integrates conditional diffusion models (CDM) with multi-task learning, enabling end-to-end joint optimization of image synthesis and anatomical segmentation. Unlike previous multitask diffusion models, our approach requires no initial predictions (e.g., a coarse segmentation mask), shares both encoder and decoder parameters across tasks, and employs a semi-supervised training strategy to learn from scans with missing segmentation labels, a common constraint in real-world clinical data. We evaluated our method on a cohort of 264 patients, where it consistently outperformed state-of-the-art single-task and multi-stage models. For image synthesis, our model achieved a PSNR of 25.61 dB, compared to 23.80 dB from a single-task CDM. For anatomical segmentation, it improved the lumen Dice score to 0.89 from 0.87 and the challenging thrombus Dice score to 0.53 from 0.48 (nnU-Net). These segmentation enhancements led to more accurate clinical measurements, reducing the lumen diameter MAE to 4.19 mm from 5.78 mm and the thrombus area error to 33.85% from 41.45% when compared to nnU-Net. Code is publicly available at https://github.com/yuxuanou623/AortaDiff.git.
@@ 1 Introduction
Clinical assessment of vascular conditions like abdominal aortic aneurysms (AAA) traditionally relies on contrast-enhanced CT (CECT), where an intravenous iodinated contrast agent illuminates the aortic lumen to reveal abnormalities such as thrombus [1]. However, this procedure poses risks to patients with renal insufficiency or contrast allergies, involves uncomfortable needle insertions, and contributes to environmental iodine waste [2, 3, 4]. This has created a significant clinical need for an advanced method capable of performing AAA assessment solely on non-contrast CT (NCCT) scans. Such an approach is particularly challenging because thrombus and blood are nearly indistinguishable in NCCT, as shown in Figure 1 (b). Recent advances in deep learning offer a promising pathway to overcome this diagnostic challenge. An effective AI solution must therefore perform two interrelated tasks: (1) generate a realistic synthetic CECT to aid visual interpretation by radiologists, and (2) accurately segment the aortic lumen, allowing thrombus to be identified as the non-enhancing regions outside this segmented area. This dual-output approach provides radiologists with both a qualitative view of the anatomy from the synthetic image and immediate quantitative measurements of the abnormality from the segmentation mask, as illustrated in Figure 1 (a).
@@
Recent AI advancements have primarily focused on generating synthetic CECTs from NCCT scans, exploring architectures from early CNNs [5] to GANs [6, 7, 8] and Transformers [9]. However, these works are constrained by a fundamental limitation: they frame the problem as a single-task, image-to-image translation. This overlooks the clinical reality that a visually plausible image is not the end goal, but an intermediate step towards quantitative analysis. Consequently, when segmentation is required, it is relegated to a separate, downstream model [7, 8]. This multi-stage pipeline is inherently flawed, as it is susceptible to error accumulation—where synthesis artifacts degrade segmentation accuracy—and inhibits knowledge sharing between the two synergistic tasks.
@@
To address the shortcomings of multi-stage methods, some pioneering works have explored unified, multi-task frameworks. For instance, Hu et al. [10] and UCAS [11] proposed models that jointly perform CECT synthesis and segmentation. However, these early multi-task attempts are predominantly built upon GAN-based architectures. While innovative in their task formulation, they inherit the well-known challenges of GANs: they often produce overly smooth outputs that lack the sharp anatomical boundaries critical for clinical diagnosis and can suffer from training instability or mode collapse [12]. Therefore, a significant gap remains for a framework that not only adopts a unified multi-task design but also leverages a more powerful generative model to ensure both high-fidelity synthesis and precise, simultaneous segmentation.
@@
To address these limitations, we propose a unified multi-task diffusion framework (AortaDiff) that jointly performs image translation from NCCT to CECT and lumen segmentation. The core of our approach is a novel denoising U-Net architecture featuring a shared encoder-decoder backbone that ends in two different prediction heads: one for the denoised image and one for the segmentation mask. This single-backbone design allows the network to learn a rich, shared latent representation that includes both texture of contrast enhancement and precise anatomical information about lumen thrombus boundary. What’s more, our model advances beyond prior multitask diffusion methods in two ways. First, it operates directly on the raw NCCT input, eliminating the need for any coarse initial predictions (e.g., a pre-computed coarse segmentation mask) during both training and inference. Second, to address the specific challenge of scarce lumen segmentation labels, we introduce a semi-supervised training strategy. This strategy incorporates a loss function that dynamically changes between supervised objectives on fully labeled data with an unsupervised reconstruction objective on image-only pairs. This allows our model to effectively leverage the entire dataset, even when only a subset of scans have lumen masks, enhancing its generalization capabilities under limited supervision and aligning with the practical constraints of clinical data acquisition.
@@
We conducted a comprehensive evaluation of our framework on the Oxford Abdominal Aortic Aneurysm (OxAAA) dataset [13]. Our unified model was benchmarked against state-of-the-art baselines and traditional multi-stage pipelines. The results demonstrate a significant and consistent improvement across all tasks. For the challenging thrombus segmentation, our model increases the Dice score from 0.48 (nnU-Net) to 0.53, and for lumen segmentation, from 0.87 to 0.89. In the synthesis task, our framework achieves a PSNR of 25.61 dB, markedly outperforming the single-task CDM’s 23.80 dB. Qualitatively, our jointly generated outputs exhibit sharper anatomical details and superior segmentation consistency. We also validate our method on two AAA clinical measurements and found that it yields more accurate morphological predictions for both lumen diameter and thrombus area, underscoring its potential for reliable clinical application.
@@
  • We propose a novel diffusion-based multi-task framework designed to move beyond simple image synthesis towards clinically actionable analysis. Based on the insight that a synthetic CECT is an intermediate step, our model jointly performs synthetic CECT generation and lumen segmentation. This integrated approach enables a safer and more sustainable workflow for AAA imaging by reducing the need for contrast agent injection.
@@ 2 Related Work: Multi-task Diffusion Models
Extending diffusion models for multi-task learning is an active area of research, with many foundational strategies first developed in general computer vision. For instance, CoDi [14] enables ’any-to-any’ generation by aligning separate, pre-trained diffusers in a shared latent space via cross-attention. TaskDiffusion [15] introduces a joint denoising process with a cross-task decoder to unify dense predictions, while DiffusionMTL [16] uses a shared encoder that branches into task-specific decoders, conditioning the diffusion process on their initial outputs. Other works, like Shadow Diffusion [17], require a coarse initial mask to guide the joint refinement process. Focusing on efficiency, MTU [18] adapts pre-trained models by replacing network layers with smaller, task-specific feed-forward network and proposing a routing mechanism. While these general-purpose frameworks are foundational, they are not tailored for the unique constraints of medical imaging, such as the need for high anatomical precision and handling of scarce clinical data.
@@
Within the medical domain, a few works have begun to adapt these concepts. DiffAtlas [19], for example, applied a joint image-mask generation model to medical images but was limited by its requirement for full supervision. While all these works have significantly advanced multi-task diffusion, a critical gap remains for a framework that is simultaneously (1) architecturally unified and specialized for a spatially coherent medical task, (2) initialization-free and label-efficient via a semi-supervised scheme, and (3) validated on a high-stakes clinical application. Our work is designed to fill this specific void.
3 Proposed Method: AortaDiff
We propose AortaDiff, a novel multi-task diffusion-based framework that jointly generates synthetic CECT images and performs lumen segmentation directly from NCCT scans. The method focuses explicitly on the aortic region, simultaneously synthesizing the CECT image and its corresponding lumen mask while preserving the surrounding background. Furthermore, we employ a semi-supervised training strategy that effectively learns both tasks by leveraging a combination of data with and without lumen mask annotations.
Figure 2: Misalignments in Paired Scans from the OxAAA Dataset. An example of spatial discrepancies between corresponding NCCT and CECT slices from the OxAAA dataset. This illustrates the inherent difficulty of achieving perfect pixel-wise alignment in practice, even after standard registration.
3.1 Aorta-focused Generation with Background Preservation
A primary challenge in paired NCCT-CECT datasets is spatial misalignment due to factors like respiratory motion, even after standard image registration [12]. An example is shown in Figure 2. To prevent the model from learning spurious translations of surrounding tissues, such as the airways, we constrain the generative process to focus exclusively on the aortic region, shown in Figure 1 (b), while preserving the original NCCT background.
Aorta-Focused Training Strategy
Inspired by [20], our training objective is localized to a bounding box tightly enclosing the aorta, which is determined from the ground truth aorta segmentation masks. Within this bounding box, we replace the standard pixel-wise MSE loss with a perceptual loss $\mathcal{L}_{p}$ [21]. This loss is less sensitive to the minor pixel shifts inherent in misaligned data and encourages the model to reconstruct the aorta texture and morphology faithfully. The reconstruction loss is therefore calculated as:
$\mathcal{L}_{\text{recon}}=\mathcal{L}_{p}(\mathbf{x}_{0}[\mathbb{B}],\mathbf{\hat{x}}_{0}[\mathbb{B}]),$ (1)
where $\mathbf{x}_{0}$ and $\mathbf{\hat{x}}_{0}$ are the ground truth and predicted CECT images, respectively, and $\mathbb{B}$ denotes the patch defined by the aorta-enclosing bounding box.
Inference and Fusion Strategy
During inference, we first obtain an aorta segmentation mask on NCCT scans using a pre-trained nnU-Net. It is trained exclusively on the training split with strict patient-level separation, using only NCCT scans and aorta labels. The model is frozen before any test evaluation and achieves a Dice score of 0.93 on the validation set. To generate the final image while preserving the background, we adapt the inpainting strategy from RePaint [22]. At each denoising timestep t, the known background region is directly replaced by the corresponding pixels from the input NCCT scan after applying the forward diffusion process noise scheduled for that step. Concurrently, the foreground aorta region is filled with our model’s denoised prediction. This composition ensures a seamless fusion of the synthesized aorta into its original anatomical context. To further enhance boundary coherence, we employ a resampling scheme [22], where the denoising process is interleaved with several noising steps, allowing the model to better harmonize the foreground and background.
3.2 A Unified Architecture for Multitask Denoising
To enable multi-task learning, we adapt the standard denoising U-Net architecture to simultaneously predict both the denoised image and its corresponding lumen segmentation. We implement this by incorporating two parallel prediction heads that branch from the final layer of the shared U-Net decoder. The first head outputs the predicted image noise, as is standard in DDPMs. For the second head, to encourage smoother and more precise lumen boundary prediction, we represent the lumen mask as a Signed Distance Function (SDF). This head therefore outputs a predicted SDF, $\mathbf{\hat{s}}_{0}$. This unified design forces the shared backbone to learn a rich latent representation beneficial for both tasks. To balance these distinct objectives, we employ the principled uncertainty-weighting strategy from Kendall et al. [23]. This approach introduces two trainable parameters, $\sigma_{1}$ and $\sigma_{2}$, to automatically learn the task weightings by optimizing the joint objective:
$\mathcal{L}(W,\sigma_{1},\sigma_{2})=\frac{1}{2\sigma_{1}^{2}}\,\mathcal{L}_{\text{seg}}(W)+\frac{1}{2\sigma_{2}^{2}}\,\mathcal{L}_{\text{recon}}(W)+\log\!\bigl(\sigma_{1}\sigma_{2}\bigr),$ (2)
where $W$ are the model parameters, $\mathcal{L}_{\text{recon}}$ is the perceptual loss on the image patches, and $\mathcal{L}_{\text{seg}}$ is the L1 loss between the predicted and ground truth SDFs: $\mathcal{L}_{\text{seg}}=\|\mathbf{\hat{s}}_{0}-\mathbf{s}_{0}\|_{1}$. This formulation provides a robust and well-founded approach to jointly optimizing both the image synthesis and geometric segmentation tasks.
3.3 Semi-Supervised Training Strategy
Figure 3: Overview of the AortaDiff Training and Inference Workflow. The model is trained using a semi-supervised strategy. For all data, the model is updated with a reconstruction loss. An additional segmentation loss is applied only when a lumen mask is available. In inference, starting with an NCCT scan and a pure Gaussian noise image, the model performs iterative denoising to jointly generate a synthetic CECT and its corresponding lumen mask.
To effectively leverage the entire OxAAA dataset, which contains 65 cases with both NCCT-CECT pair and lumen mask and 199 cases without lumen mask, we introduce a semi-supervised learning framework. Rather than incorporating complex SSL frameworks, we adopt a direct semi-supervised strategy, leveraging the inherent representation learning strengths of the diffusion model. Our strategy is designed to integrate both data types into the training process seamlessly, addressing the common challenge of label scarcity in clinical applications. The training proceeds in two phases. First, we perform a pre-training phase where the model is warmed up exclusively on the 65 fully-labeled samples for one epoch. This ensures the segmentation head receives a stable initial gradient signal. Following this, we begin the mixed-training phase, where each training batch is constructed by sampling from both the labeled and unlabeled pools at a 1:1 ratio. For samples with ground truth segmentation, the model is optimized using the combined loss mentioned in the previous subsection. For samples lacking a segmentation mask, the segmentation loss is omitted, and the network is updated solely by the reconstruction loss. This allows the shared feature backbone of our diffusion model to learn robust anatomical representations from the entire corpus of 264 cases. The overall semi-supervised training strategy is illustrated in Figure 3 (a).
AortaDiff Workflow
During training, the AortaDiff model follows the conditional diffusion model framework: the clean CECT image $\mathbf{x}_{0}$ is progressively noised to obtain a corrupted version $\mathbf{x}_{t}$, while the corresponding NCCT scan serves as the conditional input throughout the reverse denoising process. At each timestep $t$, the model receives the concatenated pair $(\mathbf{x}_{t},\text{NCCT})$ along with the timestep embedding, and predicts both the noise $\hat{\boldsymbol{\epsilon}}_{t}$ for image reconstruction and the signed distance function (SDF) $\hat{\mathbf{s}}_{0}$ for lumen segmentation. These two outputs are supervised based on label availability: when both CECT and lumen annotations are available, the model is trained using a multi-task objective. When the lumen mask is unavailable, only the reconstruction loss is applied. During inference, the model takes only the NCCT image as input. Starting from Gaussian noise, it performs iterative denoising conditioned on the NCCT over the full diffusion trajectory. The final outputs are the synthesized CECT image and the corresponding lumen mask. The inference workflow is shown in Figure 3 (b).
@@ 4 Experiments
We train and evaluate the proposed multi-tasking diffusion models, AortaDiff, on the Abdominal Aortic Aneurysm dataset released by the Oxford Abdominal Aortic Aneurysm (OxAAA) Study [8, 13]. We assess two variants of our model: AortaDiff-F, trained with full supervision on fully-labeled data, and AortaDiff-P, trained with incomplete labels using our semi-supervised strategy on the entire dataset.
@@ 4.1 OxAAA Dataset
The OxAAA dataset consists of 264 patient cases, each including spatially registered NCCT and CECT scans [13]. Aorta segmentation labels are available for all cases, while only 65 cases additionally include expert-annotated lumen masks. We use 45 of the labeled cases for training and validation, and reserve 20 for testing. In our semi-supervised setting, the remaining 199 cases without lumen annotations are used during training.
@@ 4.3 Experimental Settings
We adopt the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and no weight decay [24]. The model is trained with a batch size of 16 and an input image resolution of $256\times 256$. Training is conducted for a total of $10^{6}$ iterations.
@@ 4.4 Evaluation Metrics
For the image generation task, we evaluate the quality of the synthesized CECT slices using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [25], and Learned Perceptual Image Patch Similarity (LPIPS) [21]. Given the anatomical misalignment between the NCCT and CECT, particularly outside the aorta region, we focus our evaluation on the region of interest — the aorta. To reduce the influence of background artifacts and registration errors, we crop each slice to the smallest bounding box that encloses the aorta and compute the metrics within this cropped region.

### v2

@@ Abstract
While contrast-enhanced CT (CECT) is standard for assessing abdominal aortic aneurysms (AAA) , the required iodinated contrast agents pose significant risks, including patient allergies and environmental harm. To reduce contrast agent use, methods have focused on generating synthetic CECT from non-contrast CT (NCCT) scans. However, most adopt a multi-stage pipeline that first generates images and then performs segmentation, which leads to error accumulation and fails to leverage shared information. To address this, we propose a unified framework that generates synthetic CECT images from NCCT scans while simultaneously segmenting the aortic lumen and thrombus. Our approach integrates conditional diffusion models (CDM) with multi-task learning, enabling end-to-end joint optimization of image synthesis and anatomical segmentation. Unlike previous multitask diffusion models, our approach requires no initial predictions (e.g., a coarse segmentation mask), shares both encoder and decoder parameters across tasks, and employs a semi-supervised training strategy to learn from scans with missing segmentation labels, a common constraint in clinical data. Evaluated on a cohort of 264 patients, our method consistently outperformed state-of-the-art single-task and multi-stage models. For image synthesis, it achieved a PSNR of 25.61 dB, compared to 23.80 dB from a single-task CDM. For segmentation, it improved the lumen Dice score to 0.89 from 0.87 and the challenging thrombus Dice score to 0.53 from 0.48 (nnU-Net). These segmentation enhancements led to more accurate clinical measurements, reducing the lumen diameter MAE to 4.19 mm from 5.78 mm and the thrombus area error to 33.85% from 41.45%. Code is at https://github.com/yuxuanou623/AortaDiff.
@@ 1 Introduction
Clinical assessment of vascular conditions like abdominal aortic aneurysms (AAA) traditionally relies on contrast-enhanced CT (CECT), where an intravenous iodinated contrast agent illuminates the aortic lumen to reveal abnormalities such as thrombus [napoli2004abdominal, 10944207]. However, this procedure poses risks to patients with renal insufficiency or contrast allergies, involves uncomfortable needle insertions, and contributes to environmental iodine waste [agentharm1, agentharm2, agentharm3, 10981129]. This has created a significant clinical need for an advanced method capable of performing AAA assessment solely on non-contrast CT (NCCT) scans. Such an approach is particularly challenging because thrombus and blood are nearly indistinguishable in NCCT, as shown in Figure 1 (b). Recent advances in deep learning offer a promising pathway to overcome this diagnostic challenge. An effective AI solution must therefore perform two interrelated tasks: (1) generate a realistic synthetic CECT to aid visual interpretation by radiologists, and (2) accurately segment the aortic lumen, allowing thrombus to be identified as the non-enhancing regions outside this segmented area. This dual-output approach provides radiologists with both a qualitative view of the anatomy from the synthetic image and immediate quantitative measurements of the abnormality from the segmentation mask, as illustrated in Figure 1 (a).
@@
Recent AI advancements have primarily focused on generating synthetic CECTs from NCCT scans, exploring architectures from early CNNs [santini2018synthetic] to GANs [xie2021generation, chun2022synthetic, chandrashekar2023deep] and Transformers [ristea2023cytran]. However, these works are constrained by a fundamental limitation: they frame the problem as a single-task, image-to-image translation. This overlooks the clinical reality that a visually plausible image is not the end goal, but an intermediate step towards quantitative analysis. Consequently, when segmentation is required, it is relegated to a separate, downstream model [chun2022synthetic, chandrashekar2023deep]. This multi-stage pipeline is inherently flawed, as it is susceptible to error accumulation, where synthesis artifacts degrade segmentation accuracy, and prevents knowledge sharing between the two highly relevant tasks.
@@
To address the shortcomings of multi-stage methods, some pioneering works have explored unified, multi-task frameworks. For instance, Hu et al. [hu2022aorta] and UCAS [zhu2025partial] proposed models that jointly perform CECT synthesis and segmentation. However, these early multi-task attempts are predominantly built upon GAN-based architectures. While innovative in their task formulation, they inherit the well-known challenges of GANs: they often produce overly smooth outputs that lack the sharp anatomical boundaries critical for clinical diagnosis and can suffer from training instability or mode collapse [azarfar2023applications, 10.1007/978-3-032-06329-8_12]. Therefore, a significant gap remains for a framework that not only adopts a unified multi-task design but also leverages a more powerful generative model to ensure both high-fidelity synthesis and precise segmentation.
@@
To address these limitations, we propose a unified multi-task diffusion framework, AortaDiff, that jointly performs image translation from NCCT to CECT and lumen segmentation. The core of our approach is a novel denoising U-Net architecture featuring a shared encoder-decoder backbone that ends in two different prediction heads: one for the denoised image and one for the segmentation mask. This single-backbone design allows the network to learn a rich, shared latent representation that includes both texture of contrast enhancement and precise anatomical information about lumen thrombus boundary. What’s more, our model advances beyond prior multitask diffusion methods in two ways. First, it operates directly on the raw NCCT input, eliminating the need for any coarse initial predictions (e.g., a pre-computed coarse segmentation mask) during both training and inference. Second, to address the specific challenge of scarce lumen segmentation labels, we introduce a semi-supervised training strategy. This strategy incorporates a loss function that dynamically changes between supervised objectives on fully labeled data with an unsupervised reconstruction objective on image-only pairs. This allows our model to effectively leverage the entire dataset, even when only a subset of scans have lumen masks, enhancing its generalization capabilities under limited supervision and aligning with the practical constraints of clinical data acquisition.
@@
We conducted a comprehensive evaluation of our framework on the Oxford Abdominal Aortic Aneurysm (OxAAA) dataset [oxaaa]. Our unified model was benchmarked against state-of-the-art baselines and traditional multi-stage pipelines. The results demonstrate a consistent improvement across all tasks. For the challenging thrombus segmentation, our model increases the Dice score from 0.48 (nnU-Net) to 0.53, and for lumen segmentation, from 0.87 to 0.89. In the synthesis task, our framework achieves a PSNR of 25.61 dB, markedly outperforming the single-task CDM’s 23.80 dB. Qualitatively, our jointly generated outputs exhibit sharper anatomical details and superior segmentation consistency. We also validate our method on two AAA clinical measurements and found that it yields more accurate morphological predictions for both lumen diameter and thrombus area, underscoring its potential for reliable clinical application.
@@
  • We propose a novel diffusion-based multi-task framework designed to move beyond image synthesis towards clinically actionable analysis. Based on the insight that a synthetic CECT is an intermediate step, our model jointly performs synthetic CECT generation and lumen segmentation. This approach enables a safer and more sustainable workflow for AAA imaging by reducing the need for contrast agent injection.
@@ 2 Related Work: Multi-task Diffusion Models
Extending diffusion models for multi-task learning is an active area of research, with many foundational strategies first developed in general computer vision. For instance, CoDi [codi] enables ’any-to-any’ generation by aligning separate, pre-trained diffusers in a shared latent space via cross-attention. TaskDiffusion [taskddpm] introduces a joint denoising process with a cross-task decoder to unify dense predictions, while DiffusionMTL [diffmtl] uses a shared encoder that branches into task-specific decoders, conditioning the diffusion process on their initial outputs. Other works, like Shadow Diffusion [shadowdiffusion], require a coarse initial mask to guide the joint refinement process. Focusing on efficiency, MTU [mtu] adapts pre-trained models by replacing network layers with smaller, task-specific feed-forward network and proposing a routing mechanism. While these general-purpose frameworks are foundational, they are not tailored for the unique constraints of medical imaging, such as the need for high anatomical precision and handling of scarce clinical data.
@@
Within the medical domain, a few works have begun to adapt these concepts. DiffAtlas [diffatlas], for example, applied a joint image-mask generation model to medical images but was limited by its requirement for full supervision. While all these works have significantly advanced multi-task diffusion, a critical gap remains for a framework that is simultaneously (1) architecturally unified and specialized for a medical task, (2) initialization-free and label-efficient via a semi-supervised scheme, and (3) validated on a high-stakes clinical application. Our work is designed to fill this specific void.
@@ 3 Experiments
We train and evaluate the proposed multi-tasking diffusion models, AortaDiff, on the Abdominal Aortic Aneurysm dataset released by the Oxford Abdominal Aortic Aneurysm (OxAAA) Study [chandrashekar2023deep, oxaaa]. We assess two variants of our model: AortaDiff-F, trained with full supervision on fully-labeled data, and AortaDiff-P, trained with incomplete labels using our semi-supervised strategy on the entire dataset.
@@ 3.1 OxAAA Dataset
The OxAAA dataset consists of 264 patient cases, each including spatially registered NCCT and CECT scans [oxaaa]. Aorta segmentation labels are available for all cases, while only 65 cases additionally include expert-annotated lumen masks. We use 45 of the labeled cases for training and validation, and reserve 20 for testing. In our semi-supervised setting, the remaining 199 cases without lumen annotations are used during training.
@@ 3.3 Experimental Settings
We adopt the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and no weight decay [adamw]. The model is trained with a batch size of 16 and an input image matrix size of $256\times 256$. Training is conducted for a total of $10^{6}$ iterations.
@@ 3.4 Evaluation Metrics
For the image generation task, we evaluate the quality of the synthesized CECT slices using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [ssim], and Learned Perceptual Image Patch Similarity (LPIPS) [lpips]. Given the anatomical misalignment between the NCCT and CECT, particularly outside the aorta region, we focus our evaluation on the region of interest — the aorta. To reduce the influence of background artifacts and registration errors, we crop each slice to the smallest bounding box that encloses the aorta and compute the metrics within this cropped region.